# On the interaction of dilatancy and friction in the behavior of fluid–saturated sheared granular materials: a coupled Computational Fluid Dynamics–Discrete Element Method study

### v1

@@ III Mechanism of failure
Figure 4: (a) and (b) show the dense ($\phi$ = 0.579) and loose ($\phi$ = 0.444) subaqueous models, respectively, with points A and B located within the shear zone of the dense model and points C and D within the shear zone of the loose model. Panels (c) and (d) illustrate the dense and loose subaerial models, respectively, with points E and F located within the shear zone of the dense model and points G and H within the shear zone of the loose model.
@@
Deviatoric strain provides critical insights into material deformation under shear, revealing the development of internal failure and strain localization. Comparing dense and loose models under both subaqueous and subaerial conditions illustrates how initial packing fraction and fluid presence influence strain deformation and shear resistance. Fig. 4 presents the spatial distribution of deviatoric strain at $t=0.4\,\mathrm{s}$, highlighting contrasting shear zone patterns shaped by the interplay between particle arrangement and environmental conditions. In the dense subaqueous model (Fig. 4(a)), shear localizes into relatively narrow, well-defined zones. The interstitial fluid dampens deformation by reducing interparticle friction, suppressing abrupt strain concentration and promoting a more distributed strain field Lee (2020); Petford and Koenders (2003). Points A and B are positioned at the base and middle of the shear zone, respectively, capturing regions of concentrated yet moderated strain due to fluid support.
@@
In contrast, the loose subaqueous model (Fig. 4(b)) exhibits broader, more diffuse zones of deviatoric strain, indicative of enhanced deformation. The lower packing fraction facilitates particle rearrangement and contraction, resulting in greater mobility and higher strain magnitudes. Shear is less confined, and points C and D, located at the base and center of the shear zone, lie within a more dispersed deformation field. Under dense subaerial conditions (Fig. 4(c)), deviatoric strain is sharply localized within narrow shear bands. The absence of pore fluid increases interparticle friction, leading to pronounced strain localization and intense deformation. This observation is consistent with experimental studies on granular collapse under varying fluid conditions, which show that dense packings exhibit distinct flow regimes depending on environmental saturation Bougouin and Lacaze (2018b). Points E and F mark the bottom and mid-depth of the shear zone, aligning with regions of high strain typical of dry, densely packed materials.
@@
In the loose subaerial model (Fig. 4(d)), deviatoric strain remains moderately high but is more broadly distributed than in the dense counterpart. The low initial packing fraction promotes dilation and particle motion, while the lack of fluid amplifies rearrangement without stabilizing resistance. Consequently, points G and H, analogous in location to the other cases, fall within a wider, less confined shear zone, reflecting a more progressive failure pattern and reduced localization intensity. Overall, the figure underscores how fluid presence and initial packing density modulate the spatial extent and intensity of shear zones. Subaqueous models favor distributed deformation, while subaerial conditions promote extensive localization. Loose configurations consistently produce broader deformation zones, emphasizing the central role of initial structure in governing granular shear responses.
@@
Hence, dense models exhibit pronounced strain localization within narrow shear bands, while loose models display broader and less intense shear zones. This contrast stems from lower initial packing fractions and enhanced particle rearrangement in loose configurations. In subaqueous conditions, the presence of pore fluid further smooths the strain distribution by reducing interparticle friction. These findings highlight the critical influence of both fluid presence and packing density on strain behavior and failure dynamics. To investigate the evolution of shear zones in more detail, specific points are fixed within the deformation region. These locations enable the analysis of friction, pore pressure, and porosity by tracking associated strain rates and relevant dimensionless numbers, such as the inertial and viscous numbers. Monitoring the temporal evolution of these parameters provides insight into how shear zones behave under different mechanical and hydraulic conditions Zhou and Voyiadjis (2021); Jeon, Kim, and Kwon (2024); Parez et al. (2021). The following subsection presents results on the spatial and temporal distributions of friction, pore pressure, volumetric change, and granular rheology.
@@ IV Spatial and temporal distribution of friction, pore pressure, and volume change
Friction at selected points is calculated using coarse-grained stress tensors derived from CFD-DEM simulations. Specifically, the mobilized friction coefficient is defined as the ratio of shear stress to effective normal stress, both extracted from the local stress tensor $T_{ij}$ (Eq. 15), which accounts for all interparticle contact forces within an RVE cell. To enable meaningful comparison across different regions and packing conditions, pore pressure values were normalized by the corresponding local normal stress. While normal stress was obtained from coarse-grained stress tensors, pore pressures were sampled at selected locations near the shear zone. This normalization, performed at each time step, yielded a dimensionless pore pressure ratio that facilitates assessment of effective stress evolution and fluid–solid coupling across dense and loose granular systems. The evolution of friction, normalized pore pressure, and deviatoric strain rate over time is shown in Fig. 5, providing insight into their dynamic interplay during shear. These time-series plots delineate the transition from transient to steady-state behavior and highlight features such as stick-slip motion and strain localization Jessen, Vogt, and Cudmani (2024).
@@
Figure 5: (a) Evolution of mobilized friction coefficient at points F and H, contrasting dense ($\phi$ = 0.579) and loose ($\phi$ = 0.444) subaerial models. The mobilized friction is defined as the ratio of shear to effective normal stress, both extracted from the coarse-grained stress tensor $T_{ij}$ based on interparticle contact forces within each RVE (Eq. 15). Inset velocity profiles at 0.1 s, 0.2 s, and 1.5 s illustrate distinct failure modes. (b) Friction evolution at points A and B (dense subaqueous) and C and D (loose subaqueous), using the same stress-based definition. (c) Normalized pore pressure at the same points, defined as the instantaneous pore pressure divided by effective normal stress in the RVE at each time step, yielding a dimensionless measure of effective stress and fluid–solids interaction. (d) Deviatoric strain evolution at F and H (subaerial). (e) Deviatoric strain evolution at A–D (subaqueous), showing contrasting deformation behavior. (f) Relative porosity change ($\Delta n/n_{c}$) indicating dilation in the dense model ($n_{i}$ = 0.421, $n_{f}$ = 0.465) and compaction in the loose model ($n_{i}$ = 0.556, $n_{f}$ = 0.535), with critical porosity $n_{c}$ = 0.52.
@@
Under subaerial conditions, loosely packed granular materials initially exhibit a slightly lower friction coefficient compared to the densely packed model during the onset of failure or early transient stages (Fig. 5(a), point H). As shearing progresses, the friction coefficient gradually decreases and stabilizes, reflecting a transition to steady-state behavior characterized by stick-slip motion. This evolution is driven by progressive particle rearrangement and compaction, leading to the development of force chains and increased interparticle contacts Jessen, Vogt, and Cudmani (2024). In contrast, densely packed subaerial models exhibit slightly higher friction coefficients at failure, which remain relatively constant as the system enters a steady-state regime that also displays stick-slip characteristics (Fig. 5(a), point F). The higher friction in dense packings reflects their more rigid structure, where limited particle mobility reduces the extent of rearrangement and produces a more consistent frictional response Huang, Deng, and Wang (2024). Further insight into the mechanical behavior of these configurations is provided by the evolution of shear stress with deviatoric strain, shown in Fig. 6(a). For the densely packed model (point F), shear stress varies between approximately 0.3 and 0.4 kPa as deviatoric strain ranges from 0.08 to –0.08, indicating a relatively strong and symmetric response during forward and reverse shearing. In the loosely packed case (point H), shear stress spans from about 0.28 to 0.41 kPa over a broader strain interval of 0.15 to –0.3, reflecting a more compliant response with greater particle rearrangement. These variations highlight the contrasting shear resistance and deformation characteristics between loose and dense subaerial packings.
@@
Deviatoric strain rate (Fig. 5(d)) further highlights these differences. Dense models show lower strain rates at the start, indicative of constrained particle motion, whereas loose models display elevated strain rates during the transient phase due to extensive rearrangement, which gradually declines over time Shademani, Blais, and Shakibaeinia (2021). These strain rate trends correspond with frictional behavior: high strain rates align with low friction in loose packings, while low strain rates are associated with high friction in dense packings (Fig. 5(a), Fig. 5(d)). These patterns are consistent with experimental findings that associate increased friction with greater particle angularity, broader grain-size distributions, and enhanced surface contact, features characteristic of dense, crushed granular materials Tanikawa et al. (2025).
@@
Pore pressure plays a critical role in controlling the mechanical response of granular materials under subaqueous conditions by altering the effective stress and thereby influencing deformation and failure mechanisms Beeler et al. (2000). In loosely packed materials, larger void spaces and greater compressibility facilitate a rapid buildup of pore pressure during shearing Do et al. (2023). This is characterized by a sharp initial rise in normalized pore pressure, followed by gradual stabilization as particles rearrange and fluid redistributes through the porous structure (Fig. 5(c), points C and D). In contrast, densely packed systems initially generate negative pore pressures due to limited pore space and low compressibility, which slow pressure buildup and result in a steadier, more gradual increase (Fig. 5(c), points A and B). Despite these differences, both configurations ultimately stabilize as pore pressure equilibrates.
@@
These contrasting pore pressure dynamics strongly influence the frictional response. In loosely packed systems, elevated pore pressures reduce effective normal stress, weakening interparticle contacts and promoting particle rearrangement. This results in a transiently higher friction coefficient in loose packings (Fig. 5(b), points C and D) compared to dense packings. Shear stress in these regions ranges from 0.15 to 0.25 kPa over a deviatoric strain interval of approximately 0.02 to –1.22 (Fig. 6(b), points C and D). As shearing progresses and fluid drains from the pore space, pore pressure dissipates, causing the friction coefficient to decline and eventually stabilize at a steady-state value comparable to dense subaerial conditions Niu et al. (2024).
@@
Conversely, dense subaqueous packings exhibit lower initial friction (Fig. 5(b), points A and B) due to suction generated by negative pore pressures (Fig. 5(c), points A and B) during dilative deformation . This increases effective stress and reinforces interparticle contacts, limiting particle mobility. Shear stress in these regions ranges from 0.15 to 0.3 kPa as deviatoric strain evolves from 0.02 to –1.45 (Fig. 6(b), points A and B)), reflecting a steady buildup of shear resistance and the formation of force chains. These mechanical responses are mirrored in strain localization trends. Loose packings show higher deviatoric strain rates during the early stages of deformation, driven by enhanced particle rearrangement and compaction Jessen, Vogt, and Cudmani (2024), whereas dense systems display lower strain rates due to dilative hardening and structural stability (Fig. 5(e)). Together, these observations underscore the critical influence of initial packing and pore fluid feedback in shaping both the frictional and kinematic behavior of subaqueous granular materials.
@@
The relative change in porosity ($\Delta n/n_{c}$) provides critical insight into the volumetric behavior of granular materials during shear deformation, as illustrated in Fig. 5(f). It reveals distinct trends of dilation in dense models and compaction in loose ones. In the dense model, for example, porosity increases from an initial value of $n_{i}=0.421$ to a final value of $n_{f}=0.465$, indicating dilation. This behavior arises from particle rearrangement and the formation of void spaces, as the material resists further compaction under shear. The critical porosity ($n_{c}=0.52$), derived from poromechanics principles, serves as a threshold beyond which dilation dominates the volumetric response Shi et al. (2021).
@@
Conversely, the loose model exhibits compaction, with porosity decreasing from an initial value of $n_{i}=0.556$ to a final value of $n_{f}=0.535$, as shown in Fig. 5(f). This reduction results from the collapse of void spaces and an increase in particle contacts under shear stress, leading to a denser and more stable configuration. Comparing these behaviors highlights the influence of initial packing density on dilation and compaction mechanisms Shi et al. (2021). Dense materials resist further compression and tend to dilate, whereas loose materials undergo compaction as they transition toward a more tightly packed state Kruyt and Rothenburg (2016). These trends underscore the critical role of porosity evolution in governing shear localization and the overall mechanical response of granular materials Kothari and Elbanna (2017).

### v2

@@ III Mechanism of failure
Figure 5: (a) and (b) show the dense ($\phi$ = 0.579) and loose ($\phi$ = 0.444) subaqueous models, respectively, with points A and B located within the shear zone of the dense configuration, and points C and D within the shear zone of the loose configuration. Panels (c) and (d) illustrate the dense and loose subaerial models, respectively, with points E and F located within the shear zone of the dense configuration, and points G and H within the shear zone of the loose configuration.
Figure 6: Spatial distribution of normal effective stress $\sigma^{\prime}_{n}$, for subaqueous configurations at $t=0.1$, $0.2$, and $1.5$ s. Panels (a–c) correspond to the dense packing, and panels (d–f) to the loose packing.
Figure 7: Spatial distribution of viscous dissipation, $\psi$ (W m-3), for subaqueous configurations at $t=0.1$, $0.2$, and $1.5$ s. Panels (a–c) are for dense packing, and panels (d–f) for loose packing. Dissipation concentrates near the failure front and is more diffuse in the loose case. We compute, $\psi$ using $\psi=2\mu\,\mathbf{D}:\mathbf{D}-\tfrac{2}{3}\mu(\nabla\cdot\mathbf{u})^{2}$, which reduces to $2\mu\,\mathbf{D}:\mathbf{D}$, when $\nabla\cdot\mathbf{u}=0$.
In contrast, the loose subaqueous configuration experiences greater displacement, as shown in Fig. 2(c) and Fig. 2(d). The particles in this configuration are initially loosely packed, resulting in a lower packing fraction and more intergranular void space. This allows for particle rearrangement and contraction during shearing. As shearing progresses, contraction reduces the overall volume and diminishes shear strength due to the loss of effective interlocking, an effect well documented in subaqueous and analogous slope failure studies Zhang2021; Carey2019. This internal reorganization leads to more pronounced motion and a longer runout compared to the dense configuration. Experimental and numerical studies have further shown that loose granular configurations, which are highly sensitive to pore pressure, exhibit increased flow mobility under submerged conditions Lee2022. The loose subaqueous configuration also displays higher velocity, reflecting a more dynamic collapse process. Overall, the contrast between the dense and loose subaqueous configurations highlights the failure behavior governed by the critical role of initial packing density and fluid–particle interactions. While the dense configuration benefits from shear-induced dilation and associated frictional strengthening, the loose configuration, dominated by contraction, increased pore pressure, and reduced interparticle interlocking, undergoes a more mobile and extensive failure.
In subaerial systems, where fluid effects are absent, failure mechanisms are primarily governed by the initial packing fraction of the granular material. The dense configuration, illustrated in Fig. 2(e) and Fig. 2(f), demonstrates failure when internal stresses exceed the material’s strength. Although the collapse is dilative, meaning that volume increases as particles are sheared, the runout distance remains comparable to that of the loose subaerial case. This limited difference arises from the high initial packing fraction, which enhances interlocking and resistance to deformation Tapia2013. In the loose subaerial configuration, shown in Fig. 2(g) and Fig. 2(h), exhibits runout, extending to 0.15 meters. The lower initial packing fraction allows particles to rearrange more readily and contract during shearing, leading to increased mobility, which is comparable to the dense configuration. This contraction reduces shear strength and promotes more extensive displacement, as loose grains are less able to resist movement. Experimental findings confirm that such configurations lead to faster and broader collapses, emphasizing the importance of initial packing in governing failure dynamics under dry conditions bougouin2018granular. Consequently, the contrast between dense and loose subaerial configurations underscores the critical role of granular structure in determining collapse behavior in the absence of particle–fluid interactions.
III.1 Comparison of simulated and experimental runouts
The focus of our study is on investigating the effects of packing fraction and dilatancy on the failure and flow behavior of submerged and saturated granular deposits, and we do not aim to develop a calibrated model of any specific experiment. However, we make an attempt here to compare the behavior of the dense submerged model in our simulations with that observed in experiments on similar materials and dense packing fractions. Fig. 3(a) compares the dense immersed front position in our CFD-DEM model against the flume data of Rondon et al. (2011, their Fig. 3A). The CFD–DEM series ($\phi\ =0.579$, $A=0.80$, 3D, laminar) and the digitized measurements ($\phi\ =0.60$, $A=0.67$) exhibit the same rapid rise followed by a slow approach to a plateau; absolute offsets are attributable to aspect ratio, dimensionality (3D vs. quasi-2D), and sidewall effects. Fig. 3(b) reports the normalized runout $L/L_{i}$ for dense and loose subaerial and subaqueous cases, alongside panel (a) front-position comparisons to Rondon et al. (2011). At $A\approx 0.67$–0.80, Rondon et al. measured $L_{f}/L_{i}=2.05$ for subaqueous $\phi=0.60$, $3.08$ for subaqueous $\phi=0.55$, and $2.6$ for subaerial $\phi=0.60$. Our MFIX–DEM runs yield $L_{f}/L_{i}=2.92$ for the subaqueous dense case ($\phi=0.579$) and $3.21$ for the subaqueous loose case ($\phi=0.444$), both consistent with the experimental envelope for $\phi=0.55$–0.60.
In contrast, the subaerial cases give much larger values, $L_{f}/L_{i}=3.69$ for $\phi=0.579$ and $4.01$ for $\phi=0.444$. These higher values reflect differences in configuration: our fully 3D, flat-bed setup minimizes sidewall dissipation relative to quasi-2D flumes, a factor known to extend runouts in dry conditions Jop2005. The loose state in our simulations ($\phi=0.444$) lies below the packing range tested experimentally by Rondon and Polanía and densifies rapidly in the laminar regime, preventing the sustained pore-pressure buildup that can enhance immersed mobility in flumes. Gate-release protocol also matters. In our simulations, the front wall is removed instantaneously at t=0t{=}0, yielding an impulsive loss of support and a sharp initial acceleration. In quasi-2D flumes, laboratory collapses are typically initiated by withdrawing a swinging gate or a sliding door with a finite opening time, which transiently supports the column and modifies early shear and pore-pressure transients; this alters the impulse delivered to the front and can shift $L_{f}/L_{i}$ at fixed $(\phi,A)$ Balmforth2005; Rondon2011. Documented dry-column experiments explicitly note such gate mechanisms (“withdrawing a swinging gate or sliding door”), reinforcing that actuation trajectory and speed are first-order controls on initial momentum and, hence, on runout Balmforth2005. Nevertheless, the fundamental ordering (subaerial >>  subaqueous) is preserved across both simulations and experiments, consistent with the physics of dense column releases.
III.2 Shear-zone and pore-pressure contrasts in dense vs. loose packings
There are contrasting behaviors between dense and loose subaqueous configurations in terms of shear zone evolution and pore pressure dynamics. Dense specimens exhibit a slower, progressive build-up of pore pressure under undrained cyclic loading, attributed to their higher resistance to deformation and reduced compressibility at greater densities Bai2023. In contrast, loose configurations show broader, more diffuse shear zones accompanied by decreasing pore pressure, as fluid redistributes during contraction and particle rearrangement. This behavior is well captured in multiphase modeling studies, which demonstrate that shear-induced volume changes in loosely packed columns lead to distinct pore-pressure feedback mechanisms, influencing both the timing and extent of collapse Lee2020. Experimental studies further underscore the role of fluid: under saturated conditions, reduced flow mobility and increased resistance promote denser frontal accumulation Tong2022. Fig. 4 provides a detailed visualization of these behaviors at two distinct time steps, 0.1 seconds and 0.8 seconds, highlighting the influence of initial packing density on deformation patterns and pore pressure evolution in fluid-saturated granular columns.
Shear zones in the dense subaqueous configuration correspond to localized regions of deformation where volumetric strain approaches zero, an interpretation consistent with the critical state framework Wood1991. At the early time step of 0.1 seconds, Fig. 4(a), these zones initiate near the base and extend upward in a concentrated pattern. By 0.8 seconds, Fig. 4(b), the shear zones become more pronounced and well-defined, reflecting the progressive nature of shearing during failure. The high initial packing fraction in dense granular configurations enhances frictional resistance, limits deformation, and promotes the development of narrower, more stable shear zones over time Bai2023; Desrues2004. These features align with models of dilative deformation in dense, low-permeability granular media, where pore pressure increases due to particle interlocking and restricted fluid redistribution Petford2003. Initially, pore pressure remains low in the dense subaqueous configuration because the tightly packed structure inhibits fluid mobility, as shown in Fig. 4(c). As shearing progresses, Fig. 4(d), pore pressure gradually increases, reaching higher values by 0.8 seconds. This rise results from excess pressure generation driven by shearing-induced dilation and the limited ability of fluids to dissipate, characteristics observed in compacted, low-permeability sediments in subduction zones Nikolinakou2023. The panels illustrate how pore pressure intensifies in tandem with the development and localization of shear zones.
The shear zones in the loose subaqueous configuration exhibit broader and more diffuse deformation patterns compared to those in the dense configuration. At 0.1 seconds, Fig. 4(e), these zones appear wider and less sharply defined, reflecting the lower initial packing fraction, which promotes particle rearrangement and volumetric contraction. By 0.8 seconds, Fig. 4(f), the deformation zones expand further, illustrating the reduced shear resistance and enhanced mobility in loosely packed configurations. This behavior is consistent with two-phase flow simulations, which show that low-density granular columns develop widespread shear zones early during collapse, accompanied by elevated pore pressures that dissipate as the material contracts Lee2020. In such systems, shear bands propagate more readily due to higher initial pore pressures and lower effective stresses. As deformation progresses, these broad zones continue to widen, underscoring the role of initial packing fraction in controlling the kinematic spread of failure.
In the loose subaqueous configuration, pore pressure initially reaches elevated levels, Fig. 4(g), because the open grain framework allows greater fluid saturation. As shearing advances, Fig. 4(h), pore pressure decreases due to fluid drainage facilitated by contraction and increased permeability behavior that aligns with analytical models describing pressure evolution in contractant granular media Petford2003. This reduction in pore pressure supports ongoing deformation and further expansion of the shear zones by illustrating how pore pressure diminishes over time as particles reorganize and the granular structure evolves under submerged conditions.
@@ III.1 Comparison of simulated and experimental runouts
Deviatoric strain provides critical insights into material deformation under shear, revealing the development of internal failure and strain localization. Comparing dense and loose configurations under both subaqueous and subaerial conditions illustrates how initial packing fraction and fluid presence influence strain deformation and shear resistance. Fig. 5 presents the spatial distribution of deviatoric strain at $t=0.4\,\mathrm{s}$, highlighting contrasting shear zone patterns shaped by the interplay between particle arrangement and environmental conditions. In the dense subaqueous configuration (Fig. 5(a)), shear localizes into relatively narrow, well-defined zones. The interstitial fluid dampens deformation by reducing interparticle friction, suppressing abrupt strain concentration and promoting a more distributed strain field Lee2020; Petford2003. Points A and B are positioned at the base and middle of the shear zone, respectively, capturing regions of concentrated yet moderated strain due to fluid support.
@@
In contrast, the loose subaqueous configuration (Fig. 5(b)) exhibits broader, more diffuse zones of deviatoric strain, indicative of enhanced deformation. The lower packing fraction facilitates particle rearrangement and contraction, resulting in greater mobility and higher strain magnitudes. Shear is less confined, and points C and D, located at the base and center of the shear zone, lie within a more dispersed deformation field. Under dense subaerial conditions (Fig. 5(c)), deviatoric strain is sharply localized within narrow shear bands. The absence of pore fluid increases interparticle friction, leading to pronounced strain localization and intense deformation. This observation is consistent with experimental studies on granular collapse under varying fluid conditions, which show that dense packings exhibit distinct flow regimes depending on environmental saturation bougouin2018granular. Points E and F mark the bottom and mid-depth of the shear zone, aligning with regions of high strain typical of dry, densely packed materials.
@@
In the loose subaerial configuration (Fig. 5(d)), deviatoric strain remains moderately high but is more broadly distributed than in the dense counterpart. The low initial packing fraction promotes dilation and particle motion, while the lack of fluid amplifies rearrangement without stabilizing resistance. Consequently, points G and H, analogous in location to the other cases, fall within a wider, less confined shear zone, reflecting a more progressive failure pattern and reduced localization intensity. Overall, the figure underscores how fluid presence and initial packing density modulate the spatial extent and intensity of shear zones. Subaqueous models favor distributed deformation, while subaerial conditions promote extensive localization. Loose configurations consistently produce broader deformation zones, emphasizing the central role of initial structure in governing granular shear responses.
@@
We further observe that the dense configurations exhibit pronounced strain localization within narrow shear bands, while loose configurations display broader and less intense shear zones. This contrast stems from lower initial packing fractions and enhanced particle rearrangement in loose configurations. To investigate the evolution of shear zones in more detail, specific points are fixed within the deformation region. These locations enable the analysis of friction, pore pressure, and porosity by tracking associated strain rates and relevant dimensionless numbers, such as the inertial and viscous numbers. Monitoring the temporal evolution of these parameters provides insight into how shear zones behave under different loading conditions. Section IV presents results on the spatial and temporal distributions of friction, pore pressure, volumetric change, and granular rheology at these points in the shear zone. .
Figures 6 and 7 provide additional diagnostics of ambient–fluid effects. Figure 6 shows that the effective normal stress, $\sigma^{\prime}_{n}\equiv p^{\prime}$, remains strictly positive across the shear band for both dense and loose packings at all sampled times. The dense configuration largely preserves its initial value for a finite period after failure initiation (gate opening). By contrast, the loose configuration exhibits a pronounced reduction in effective normal stress and a higher initial fluidization potential shortly after gate opening. Figure 7 shows that viscous dissipation, $\psi$, localizes near the failure front, the area with the highest granular rearrangements, and is more diffuse in the loose case.
@@ IV Spatial and temporal distribution of friction, pore pressure, and volume change
Friction at selected points is calculated using coarse-grained stress tensors derived from CFD-DEM simulations. Specifically, the mobilized friction coefficient is defined as the ratio of shear stress to effective normal stress, both extracted from the local stress tensor $T_{ij}$ (Eq. 17), which accounts for all interparticle contact forces within an RVE cell. To enable meaningful comparison across different regions and packing conditions, pore pressure values were normalized by the corresponding local normal stress. While normal stress was obtained from coarse-grained stress tensors, pore pressures were sampled at selected locations near the shear zone. This normalization, performed at each time step, yielded a dimensionless pore pressure ratio that facilitates assessment of effective stress evolution and fluid–solid coupling across dense and loose granular systems. The evolution of friction, normalized pore pressure, and deviatoric strain rate over time is shown in Fig. 8, providing insight into their dynamic interplay during shear. These time-series plots delineate the transition from transient to steady-state behavior and highlight features such as stick-slip motion and strain localization Jessen2024. We track the time series of the apparent friction, $\mu_{\mathrm{}}(t)=\tau/\sigma^{\prime}_{n}(t)$, and the normalized pore-pressure ratio, $p(t)/\sigma^{\prime}_{n}(t)$, at labeled shear-band points to quantify transient and relaxation prior to steady state.
@@
Figure 8: (a) Evolution of mobilized friction coefficient at points F and H, contrasting dense ($\phi$ = 0.579) and loose ($\phi$ = 0.444) subaerial configurations. The mobilized friction is defined as the ratio of shear to effective normal stress, both extracted from the coarse-grained stress tensor $T_{ij}$ based on interparticle contact forces within each RVE (Eq. 17). Inset velocity profiles at 0.1 s, 0.2 s, and 1.5 s illustrate distinct failure modes. (b) Friction evolution at points A and B (dense subaqueous) and C and D (loose subaqueous), using the same stress-based definition. (c) Normalized pore pressure at the same points, defined as the instantaneous pore pressure divided by effective normal stress in the RVE at each time step, yielding a dimensionless measure of effective stress and fluid–solids interaction. (d) Deviatoric strain evolution at F and H (subaerial). (e) Deviatoric strain evolution at A–D (subaqueous), showing contrasting deformation behavior. (f) Relative porosity change ($\Delta n/n_{c}$) indicating dilation in the dense ($n_{i}$ = 0.421, $n_{f}$ = 0.465) and compaction in the loose ($n_{i}$ = 0.556, $n_{f}$ = 0.535), with critical porosity $n_{c}$ = 0.52.
@@
Under subaerial conditions, loosely packed granular materials initially exhibit a slightly lower friction coefficient compared to the densely packed model during the onset of failure or early transient stages (Fig. 8(a), point H). As shearing progresses, the friction coefficient gradually decreases and stabilizes, reflecting a transition to steady-state behavior characterized by stick-slip motion. This evolution is driven by progressive particle rearrangement and compaction, leading to the development of force chains and increased interparticle contacts Jessen2024. In contrast, densely packed subaerial models exhibit slightly higher friction coefficients at failure, which remain relatively constant as the system enters a steady-state regime that also displays stick-slip characteristics (Fig. 8(a), point F). The higher friction in dense packings reflects their more rigid structure, where limited particle mobility reduces the extent of rearrangement and produces a more consistent frictional response Huang2024. Further insight into the mechanical behavior of these configurations is provided by the evolution of shear stress with deviatoric strain, shown in Fig. 9(a). For the densely packed model (point F), shear stress varies between approximately 0.3 and 0.4 kPa as deviatoric strain ranges from 0.08 to –0.08, indicating a relatively strong and symmetric response during forward and reverse shearing. In the loosely packed case (point H), shear stress spans from about 0.28 to 0.41 kPa over a broader strain interval of 0.15 to –0.3, reflecting a more compliant response with greater particle rearrangement. These variations highlight the contrasting shear resistance and deformation characteristics between loose and dense subaerial packings.
@@
Deviatoric strain rate (Fig. 8(d)) further highlights these differences. Dense configurations show lower strain rates at the start, indicative of constrained particle motion, whereas loose configurations display elevated strain rates during the transient phase due to extensive rearrangement, which gradually declines over time Shademani2021. These strain rate trends correspond with frictional behavior: high strain rates align with low friction in loose packings, while low strain rates are associated with high friction in dense packings (Fig. 8(a), Fig. 8(d)). These patterns are consistent with experimental findings that associate increased friction with greater particle angularity, broader grain-size distributions, and enhanced surface contact, features characteristic of dense, crushed granular materials Tanikawa2025.
@@
Pore pressure plays a critical role in controlling the mechanical response of granular materials under subaqueous conditions by altering the effective stress and thereby influencing deformation and failure mechanisms Beeler2000. In loosely packed materials, larger void spaces and greater compressibility facilitate a rapid buildup of pore pressure during shearing Do2023. This is characterized by a sharp initial rise in normalized pore pressure, followed by gradual stabilization as particles rearrange and fluid redistributes through the porous structure (Fig. 8(c), points C and D). In contrast, densely packed systems initially generate negative pore pressures due to limited pore space and low compressibility, which slows pressure buildup and results in a steadier, more gradual increase (Fig. 8(c), points A and B). Despite these differences, both configurations ultimately stabilize as pore pressure equilibrates.
@@
These contrasting pore pressure dynamics strongly influence the frictional response. In loosely packed systems, elevated pore pressures reduce effective normal stress, weakening interparticle contacts and promoting particle rearrangement. This results in a transiently higher friction coefficient in loose packings (Fig. 8(b), points C and D) compared to dense packings. Shear stress in these regions ranges from 0.15 to 0.25 kPa over a deviatoric strain interval of approximately 0.02 to –1.22 (Fig. 9(b), points C and D). As shearing progresses and fluid drains from the pore space, pore pressure dissipates, causing the friction coefficient to decline and eventually stabilize at a steady-state value comparable to dense subaerial conditions Niu2024.
@@
Conversely, dense subaqueous packings exhibit lower initial friction (Fig. 8(b), points A and B) due to suction generated by negative pore pressures (Fig. 8(c), points A and B) during dilative deformation . This increases effective stress and reinforces interparticle contacts, limiting particle mobility. Shear stress in these regions ranges from 0.15 to 0.3 kPa as deviatoric strain evolves from 0.02 to –1.45 (Fig. 9(b), points A and B)), reflecting a steady buildup of shear resistance and the formation of force chains. These mechanical responses are mirrored in strain localization trends. Loose packings show higher deviatoric strain rates during the early stages of deformation, driven by enhanced particle rearrangement and compaction Jessen2024, whereas dense systems display lower strain rates due to dilative hardening and structural stability (Fig. 8(e)). Together, these observations underscore the critical influence of initial packing and pore fluid feedback in shaping both the frictional and kinematic behavior of subaqueous granular materials.
@@
The relative change in porosity ($\Delta n/n_{c}$) provides critical insight into the volumetric behavior of granular materials during shear deformation, as illustrated in Fig. 8(f). It reveals distinct trends of dilation in dense configurations and compaction in loose ones. In the dense configuration, for example, porosity increases from an initial value of $n_{i}=0.421$ to a final value of $n_{f}=0.465$, indicating dilation. This behavior arises from particle rearrangement and the formation of void spaces, as the material resists further compaction under shear. The critical porosity ($n_{c}=0.52$), derived from poromechanics principles, serves as a threshold beyond which dilation dominates the volumetric response Shi2021.
@@
Conversely, the loose configuration exhibits compaction, with porosity decreasing from an initial value of $n_{i}=0.556$ to a final value of $n_{f}=0.535$, as shown in Fig. 8(f). This reduction results from the collapse of void spaces and an increase in particle contacts under shear stress, leading to a denser and more stable configuration. Comparing these behaviors highlights the influence of initial packing density on dilation and compaction mechanisms Shi2021. Dense materials resist further compression and tend to dilate, whereas loose materials undergo compaction as they transition toward a more tightly packed state Kruyt2016. These trends underscore the critical role of porosity evolution in governing shear localization and the overall mechanical response of granular materials Kothari2017.